OpenStudy (underdawg215):

What is the solution of this linear system? x + 3y = -8 4x − 3y = 23

OpenStudy (anonymous):

add both equations x+3y+4x-3y=-8+23 5x=15 x=3 put x=3 in any one equation 3+3y=-8 y=-11/3

OpenStudy (underdawg215):

thank you

OpenStudy (anonymous):

ur welcome :)